When light of wavelength 120 nm falls on a gold suWhen light of wavelength 120 nm falls on a gold surface, electrons having a maximum kinetic energy of 5.25 eV are emitted. Find values for the following.
Answer #1

A photon of 120 nm will have

E = hc/lamda

E = 1249/120 = 10.4 eV

Work funtion = Energy - K max

W F = 10.4 - 5.25 = 5.15 eV

Cutoff wavelength = hc/WF = 1249 /5.15 = 242.5 nm

Frequncy = c/wavelgnth = 1.23 1015 Hz